OpenStudy (anonymous):

solve x^3-5x >= 4x^2

OpenStudy (anonymous):

(greater than or equal to)

OpenStudy (lgbasallote):

\[x^3 - 5x \ge 4x^2\] subtract 4x^2 from both sides \[\implies x^3 - 4x^2 - 5x \ge 0\] factor out x \[\implies x(x^2 - 4x - 5) \ge 0\] factor out the trinomial \[\implies x(x-5)(x+1) \ge 0\] does that help?

OpenStudy (anonymous):

I got that far, I just need to know the solutions written as integrals.

OpenStudy (anonymous):

Actually, I got it from here. Thanks!